# Updatable Balanced Index for Fast On-device Search with Auto-selection Model

### v1

@@ I Introduction
Edge devices such as autonomous vehicles [WangBCC21] and drones [Xiao24] are typically equipped with sensors like GPS and lidar. These sensors continuously generate large volumes of data points [He0KXJ0022], such as 2D geo-location from GPS and 3D point clouds [Chang2019] from lidar. The on-device search, such as kkNN search and radius search, can be applied to such data for efficient analysis and learning [XuTZ19, IP22, MatthiesHCJ08, Xu2023, Zhu22]. For example, kkNN search is the core idea behind several clustering accelerators [Yushuai2024, Yushuai2025], which can improve $\mathrm{k}$-means efficiency in classification [Lang2019] and segmentation [Qi2017, Hu2020]. 4Sheng Wang is the corresponding author.
@@
Inconsistent Query Performance.  No single search strategy can perform well for all queries. For example, in Fig. 1(b), we perform kkNN queries with two points, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$. We use two traversal strategies to explore the search space: Breadth-First Search (BFS) and Depth-First Search (DFS). The DFS-based strategy is more efficient for $\mathbf{p}_{1}$, while the BFS-based strategy is better for $\mathbf{p}_{2}$. This is because the dimensionality and distribution of the dataset have varying effects on the DFS and BFS search efficiency [SternKFH10]. Manual selection of optimal strategies is impractical due to lacking a simple and discernible rule [Wolpert1997]. A feasible solution to improve selecting optimal strategies’ efficiency is using auto-selection models [SONG2012]. However, existing auto-selection methods [Abdullah2022, Wang2020] encounter inefficiencies during training sample generation, attributed to the lack of a fast and informative feature extraction method.
@@ II Related Work
Most auto-selection models [SONG2012, Abdullah2022, Wang2020] aim to find the fastest or most accurate method for a given task. For example, Unik [Wang2020] models the fastest $\mathrm{k}$-means algorithm selection as a multi-label classification problem. It extracts features and trains a classifier from samples to predict the optimal algorithm pipeline for accelerating $\mathrm{k}$-means clustering tasks. However, Unik only extracts basic features for each task, such as the data volume for the dataset, which cannot gain a high prediction accuracy for arbitrary datasets. TRIO [Shapira2021] extracts more features to improve prediction accuracy, yet it overlooks the time costs, for instance, those incurred by utilizing the graph convolutional neural network [li2018adaptive] for feature extraction.
@@ IV-B Accelerating Pivot Set Finding
Input: $\mathbf{x}$: the training set.
@@ 5  Store the vectors in NN and return NN;
7Find $t$ by the distribution-aware partition method;
@@ VI Search Strategy Auto-selection
Output: $Q$: the bounded INqueue.
@@ VI-A Search Strategies
Numerous factors could impact the query efficiency, and considering all possible factors in the search strategies is computationally impractical. Therefore, we narrow our focus to two key factors: traversal strategies (DFS or BFS) and the minimum bounding volume (MBB or MBR), resulting in four search strategies. Table II outlines the pruning strategies corresponding to each search strategy for kkNN and radius search. For example, $R_{{{\texttt{DFS}}}}$ represents a search strategy employing the DFS and the MBR. Specifically, in kkNN search, Lemma 3 is utilized to prune search spaces, while in radius search, Lemma 1 is employed for the same purpose. Due to space limitations, we only provide the kkNN procedure here, while the detailed radius search process is included in the appendix (see Section IX-A). Specifically, in kkNN, both traversal strategies and pruning are essential for query efficiency, whereas in radius search, only traversal strategies have a notable impact.
@@ VI-B1 Class Label Generation
Input: $\mathbf{D}$: the dataset, $\mathcal{G}$: the query task set.
Output: $z^{*}$: the optimal search strategy.
1 $\mathcal{G}_{0}\leftarrow$ Randomly generated query task set;
2 for each query task $g\in\mathcal{G}_{0}$ do
3   Extracting features $\mathbf{F}$ from $g$;
4 for each search strategy $z\in\mathbf{Z}$ do
5      Running $g$ using the search strategy $z$;
6
7 $z^{*}\leftarrow$ the fast search strategy;
8 $S$.add($\mathbf{F},z^{*}$>);
@@ 7 $z^{*}\leftarrow$ the fast search strategy;
10 Training the classifier on $S$;
11 A new query task $g$;
12 $\mathbf{F}\leftarrow$ extracting features from $g$;
13 Predicting the optimal search strategy $z^{*}$ for $g$;
14 return $z^{*}$;
@@ VII Experiments
TABLE IV: Impact of parameters on UnIS construction.
Dataset Parameters Sampling Rate $\delta$ # Models $l$
@@ VII-C Index Insertion
Parameter Study.  Table VI presents the insertion latency of our index under varying parameters.
@@ VII-D Evaluation of Auto-selection Model
We first evaluate the auto-selection model and then verify it can accelerate queries. Due to space limitations, we verify its effectiveness in accelerating $\mathrm{k}$-means in the appendix (see Section IX-E).
@@
Verification for Feature Extraction.  We use the mean reciprocal rank ($MRR$) [Wang2020] to assess the auto-selection model performance. The empirical results are presented in Table VII.
@@ IX-E Verification for $\mathrm{k}$-means
Observations. As shown in Fig. 14, when $k$ is small (e.g., $k=10$), UnIS performs better than most SOTAs in pruning in most cases, but it’s not always the best. For example, Elkan outperforms it because UnIS requires additional time to construct our indexes, while $\mathtt{k}$NN on these indexes is inefficient when $k$ is small. Whereas when $k$ is large, UnIS demonstrates superior runtime performance due to the effective pruning power of the centroid index.

### v2

@@ Abstract
Index Terms:
$k$-nearest neighbor search, radius search, balanced multi-way KD-tree
@@ I Introduction
Edge devices such as autonomous vehicles [WangBCC21] and drones [Xiao24] are typically equipped with sensors like GPS and lidar. These sensors continuously generate large volumes of data points, such as 2D geo-location from GPS and 3D point clouds [Chang2019] from lidar. The on-device search, such as kkNN search and radius search, can be applied to such data for efficient analysis and learning [IP22, MatthiesHCJ08, Xu2023, Zhu22]. For example, kkNN search is the core idea behind several clustering accelerators [Yushuai2024, Yushuai2025], which can improve $\mathrm{k}$-means efficiency. 1Sheng Wang is the corresponding author.
@@
Inconsistent Query Performance.  No single search strategy can perform well for all queries. For example, in Fig. 1(b), we perform kkNN queries with two points, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$. We use two traversal strategies to explore the search space: Breadth-First Search (BFS) and Depth-First Search (DFS). The DFS-based strategy is more efficient for $\mathbf{p}_{1}$, while the BFS-based strategy is better for $\mathbf{p}_{2}$. This is because the dimensionality and distribution of the dataset have varying effects on the DFS and BFS search efficiency [SternKFH10]. Manual selection of optimal strategies is impractical due to the lack of a simple and discernible rule [Wolpert1997]. A feasible solution to improve the efficiency of selecting optimal strategies is using auto-selection models [SONG2012]. However, existing models [Abdullah2022, Wang2020] encounter inefficiencies during sample generation, due to the absence of a fast and informative feature extraction method.
@@ II Related Work
Most auto-selection models [SONG2012, Abdullah2022, Wang2020] aim to find the fastest or most accurate method for a given task. For example, Unik [Wang2020] models the fastest $\mathrm{k}$-means algorithm selection as a multi-label classification problem. It extracts features and trains a classifier from samples to predict the optimal algorithm pipeline for accelerating $\mathrm{k}$-means clustering tasks. However, Unik only extracts basic features for each task, such as the data volume for the dataset, which cannot achieve a high prediction accuracy for arbitrary datasets. TRIO [Shapira2021] extracts more features to improve prediction accuracy, yet it overlooks the time costs, for instance, those incurred by utilizing the graph convolutional neural network [li2018adaptive] for feature extraction.
@@ IV-B Accelerating Pivot Set Finding
Input: $\mathbf{x}^{\prime}_{n_{1}}$: the training set.
Output: $\mathbf{M}$: the CDF model.
@@ 5  Store the vectors in NN and return NN;
7Determine $t$ using the AEPL-optimal criterion;
@@ VI Search Strategy Auto-selection
Output: $Q$: the priority queue.
@@ VI-A Search Strategies
Numerous factors could impact the query efficiency, and considering all possible factors in the search strategies is computationally impractical. Therefore, we narrow our focus to two key factors: traversal strategies (DFS or BFS) and the minimum bounding volume (MBB or MBR), resulting in four search strategies. Table II outlines the pruning strategies corresponding to each search strategy for kkNN and radius search. For example, $R_{{{\texttt{DFS}}}}$ represents a search strategy employing the DFS and the MBR. Specifically, in kkNN search, Lemma 3 is utilized to prune search spaces, while in radius search, Lemma 1 is employed for the same purpose. Due to space limitations, we only provide the kkNN procedure here, while the detailed radius search process is included in the appendix (see Section IX-A). In kkNN, both traversal strategies and pruning are essential for query efficiency, whereas in radius search, only traversal strategies have a notable impact.
@@ VI-B1 Class Label Generation
Input: $\mathbf{D}$: the dataset; $\mathcal{G}$: the query task set.
Output: $\mathcal{Z}$: the set of predicted search strategies.
1
[…]
3 $S\leftarrow$ an empty training set, $Z\leftarrow$ an empty strategy set;
4
5for each query task $g\in\mathcal{G}_{0}$ do
6   Extract features $\mathbf{F}$ from $g$;
7 $z^{*}\leftarrow$ the fastest among the four search strategies;
8 $S.\text{add}(\langle\mathbf{F},z^{*}\rangle)$;
@@ 7 $z^{*}\leftarrow$ the fastest among the four search strategies;
10
11Train the classifier on $S$;
12
13for each query task $g\in\mathcal{G}$ do
14   Extract features $\mathbf{F}$ from $g$;
15   Predict the optimal strategy $z^{*}$ for $g$ based on $\mathbf{F}$;
16 $\mathcal{Z}.\text{add}(z^{*})$;
17
18return $\mathcal{Z}$;
19
@@ VII Experiments
TABLE IV: Impact of parameter settings on UnIS construction evaluated on the ArgoPOI dataset.
Parameters Sampling Rate $\delta$ # Models $l$
@@ VII-C Index Insertion
Parameter Study.  Table VI presents the insertion latency of our index under various parameter settings. Due to page constraints, the experiments are conducted on the ArgoPOI, ArgoAVL, and Porto datasets.
@@ VII-D Evaluation of Auto-selection Model
We first evaluate the auto-selection model and then verify that it can accelerate queries. We verify its effectiveness in accelerating $\mathrm{k}$-means in the appendix (see Section IX-E).
@@
Verification for Feature Extraction.  We evaluate the performance of the auto-selection model using mean reciprocal rank ($MRR$) [Wang2020], macro-averaged precision ($\mathcal{P}_{1}$), and micro-averaged precision ($\mathcal{P}_{2}$). The results are reported in Table VII.
@@ Acknowledgments
AI-Generated Content Acknowledgement
We used ChatGPT only to check grammar issues in the manuscript. All technical ideas, algorithmic designs, experimental analyses, manuscript writing, and code were entirely completed by the authors.
@@ IX-E Verification for $\mathrm{k}$-means
Observations. As shown in Fig. 14, when $\mathrm{k}$ is small (e.g., $\mathrm{k}=10$), UnIS performs better than most SOTAs in pruning in most cases, but it’s not always the best. For example, Elkan outperforms it because UnIS requires additional time to construct our indexes, while kkNN on these indexes is inefficient when $\mathrm{k}$ is small. Whereas when $\mathrm{k}$ is large, UnIS demonstrates superior runtime performance due to the effective pruning power of the centroid index.